# Can Carbon-Aware Electric Load Shifting Reduce Emissions? An Equilibrium-Based Analysis

### v1

@@ II-B Three-bus Example Illustration
In order to illustrate the outcomes of using an equilibrium model with carbon-dependent utility for consumers, we employ a simple three-bus system adapted from Example 6.2.2 in [20], as seen in Fig. 2. The network includes three generators and three consumers, with one generator and consumer located at each bus. Since the carbon intensity of the generators is not included in the original three-bus system data, we define two carbon emission intensity cases to see different results of load shifting among consumers in Table I. In Case I, we assume that higher cost generators have lower carbon emissions. Specifically, the most expensive generator, located at Bus 2, has the lowest carbon intensity, while the cheapest generator, located at Bus 3, has the highest carbon intensity. The medium cost generator at Bus 1 has an intermediate carbon intensity. In Case II, we assume that higher cost generators have higher carbon emissions. We assume that all consumers have a carbon cost of $c_{D,d}=\$20/$tons.
TABLE I: Parameters of Generators and Consumers.
@@
(a) Case I. Higher cost generators have lower carbon intensity.
(b) Case II. Higher cost generators have higher carbon intensity.
Figure 1: Results for different generation cost-carbon intensity pairs.
(a) Case I.
(b) Case II.
Figure 2: Generation and consumption for Case I and Case II in Table I.
II-B1 Electricity Pool Case
@@
We first consider the electricity pool case, which omits consideration of transmission constraints. The results for this case are shown in Fig. 1 (a) and (b) and detailed generation dispatch (purple numbers) and demand values (green numbers) are shown in Fig. 2. Fig. 1 (a) shows results for Case I, with the black line showing the generation cost curve (with generators ordered from the cheapest to the most expensive) and the purple line showing the total carbon emission curve for Case I. The total carbon emissions $E_{tot}$ is calculated by multiplying the amount of generation with the corresponding carbon intensities, i.e., $E_{tot}=\sum_{g\in\mathcal{G}}e_{G,g}\cdot P_{G,g}$ based on the generation dispatch order. The red and purple points on the black generation cost curve represent the dispatch results for two cases, respectively. The purple shadowed areas highlight the effects of our proposed model compared with the case without considering carbon emissions. Similarly, Fig. 1 (b) shows results for Case II, with the black line showing the generation cost curve (which is the same as for Case I) and the green line showing the total emission curve for Case II.
@@
Our first observation from Fig. 1, is that the only impact of considering the carbon-dependent utility is that the consumers are less willing to pay for electricity with a non-zero average carbon footprint, and hence reduce their electricity demand. The generation merit order (represented by the generator cost curve) creates an ordering of the generators. The demand reduction sequentially impacts the generators, starting with the marginal one which reduces production, and if needed, continues in descending order. If the affected generators have non-zero emissions, reductions in demand have the intended effect of decreasing emissions, as witnessed in both Case I and Case II. However, the fact that carbon-aware loads can only sequentially impact the generators according to the aforementioned ordering, which is not related to the ranking of generators by their emissions, is a fundamental limit of carbon-aware load shifting in our equilibrium model.
@@
Our second observation from Fig. 1 is that the demand reduction and associated reduction in emissions is very different between Case I and II. In Case I, the generation dispatch includes low-cost, carbon-intensive generation which leads to a high average carbon value $\lambda=0.913$. Since $\lambda$ is high, the carbon-dependent consumer utility is much lower than the normal consumer utility, leading to a large reduction in demand from 48 to 32 MW, or 33%, when considering carbon. However, since the marginal generator has low emissions, this reduction in demand only decreases emissions from 37 to 29.2 tons, or 21%. This suggests that each unit of demand reduction achieves a lower-than-average impact on emissions in Case I. Conversely, in Case II, the generation dispatch includes low-cost, low-carbon generation which leads to a low average carbon value $\lambda=0.4$. With a lower $\lambda$, the carbon-dependent consumer utility is closer to the normal consumer utility, and the demand is only reduced to 46.7 MW, or less than 3%, when considering carbon. This reduction in demand reduces emissions from 20 to 18.7 tons, or 6.5%. Thus, in Case II, each unit of demand reduction achieves a higher-than-average impact on emissions, but the overall demand and carbon reductions are small.
@@
(a) Base case ($c_{D,d}=0$ for $\forall d\in\mathcal{D}$).
(b) Case I. Higher cost-lower carbon generators.
(c) Case II. Higher cost-higher carbon generators.
Figure 3: Results comparison for transmission-constrained cases.
II-B2 Transmission-constrained Case
@@
To explain the impacts of transmission line constraints in our equilibrium model, we reduce the transmission capacity on the transmission line from Bus $2$ to Bus $3$ so it becomes congested. This congestion increases power generation at Bus 2 and decreases power generation at Bus 1. Fig. 3 shows the dispatch for the base case ($c_{D}=0$) and Case I and Case II ($c_{D}=20$), all with transmission line limits. For Case I, the shift of generation from generator 1 to 2 increases the average carbon emissions from $\lambda=0.608$ (for electricity pool case) to $\lambda=0.767$ in the base transmission-constrained case (Case I in Fig. 3a). Introducing carbon-sensitive loads reduces the load from 48 to 32 MW, as in the case without transmission constraints. However, due to congestion, both generator 1 and 2 are now marginal and are impacted by the demand reduction leading to a slightly higher reduction in emissions of 22.5%. Nevertheless, the average emissions increase to $\lambda=0.89$ after the load reduction. For Case II, shifting generation from generator 1 to 2 reduces the average carbon emissions to $\lambda=0.433$. This $\lambda$ is low enough to allow consumers to continue to consume their maximum load, and as a result, the generation dispatch does not change after we introduce carbon-sensitive loads. This is despite the fact that the two marginal generators in this case have higher emissions than in Case I.
@@
This example highlights that transmission constraints impact the marginal generators serving a carbon-sensitive loads, and thus impacts the effectiveness of demand reductions from carbon-sensitive loads.
III Numerical Studies
@@
This section conducts extensive numerical studies to demonstrate how equilibrium method could impact electricity market clearing. All the tests are performed on a laptop computer with an Apple M3 Pro CPU and 36GB RAM. The equilibrium problem is solved using EMP [24]. We adopt the IEEE RTS-GMLC system [25] for numerical analysis. This system has $73$ buses, $158$ generators, and $120$ lines. For generator fuel types, this system covers natural gas, oil, coal, and renewable sources (solar, wind, and hydro). The emission intensities of different generators are assigned according to the US Department of Energy [26] and are listed in Table II. For simplicity, we consider each non-zero load in the system as a consumer, leading to a total of $51$ consumers located at different buses. According to [27], carbon prices in the world-wide emission trading markets range from $10$ to $100$\$/tons. Based on these statistics, we assume that consumers will have carbon prices $c_{D}$ in the range between $[10,100]$ in the following experiments. For consumers’ demand flexibility, we use the values in the IEEE RTS-GMLC system as the maximum values and 80% of these values as the minimum values.
TABLE II: Generator carbon emission intensities.
Type Carbon Emission Intensity(tons/MWh)
Natural Gas 0.6042
Oil 0.7434
Coal 0.9606
Wind, Solar, Hydro 0
III-A Evaluation Metrics and Benchmark Formulations
@@
In our case study, we consider the following quantities when evaluating market clearing solutions:
@@
We consider three benchmark models for our analysis:
@@
  • Market Clearing with Demand Flexibility (MCDF): In this case, consumers have flexibility in their electricity consumption, but do not consider carbon costs. Note that this model is equivalent to a version of the proposed equilibrium model where all consumers set carbon costs $c_{D}=0$.
  • Sequential Method: We consider the sequential method first proposed for to represent geographical load shifting in [28] and extended to consider multiple timesteps in [8]. This method consists of three steps:
@@
    Step $3$: Finally, we solve another market clearing similar to the MCDF, representing the intra-day electricity market, with updated load profiles. We denote the average carbon emission based on these market clearing results as $\lambda_{a}$.
III-B Analysis for Different Carbon Costs
We first analyze the impact of different carbon costs on the equilibrium results, assuming that all consumers are willing to submit non-zero carbon costs to help reduce carbon emissions. We compare cases where the carbon costs of consumers are randomly drawn from a uniform distribution with ranges $[10,40]$, $[30,60]$, $[50,80]$, and $[80,100]$, respectively. Table III shows the results. The top rows represent the two carbon-agnostic formulations (MCFD and MCDF), and the average carbon cost increases as we go further down in the rows.
TABLE III: The impact of different carbon costs on generation dispatch, system emissions, and other components in equilibrium models.

### v2

@@ 2.2. Three-bus Example Illustration
In order to illustrate the outcomes of using an equilibrium model with carbon-dependent utility for consumers, we employ a simple three-bus system adapted from Example 6.2.2 in (Gabriel et al., 2012), as seen in Fig. 2(b). The network includes three generators and three consumers, with one generator and consumer located at each bus. Since the carbon intensity of the generators is not included in the original three-bus system data, we define two carbon emission intensity cases to see different results of load shifting among consumers in Table 1. In Case I, we assume that higher-cost generators have lower carbon emissions. Specifically, the most expensive generator, located at Bus 2, has the lowest carbon intensity, while the cheapest generator, located at Bus 3, has the highest carbon intensity. The medium-cost generator at Bus 1 has an intermediate carbon intensity. In Case II, we assume that higher-cost generators have higher carbon emissions. We assume that all consumers have a carbon cost of $c_{D,d}=\mathdollar 20/$tons.
Table 1. Parameters of Generators and Consumers.
@@
(a) Results for different generation cost-carbon intensity pairs.
(b) Generation and consumption for Case I and Case II in Table 1.
2.2.1. Electricity Pool Case
@@
We first consider the electricity pool case, which omits consideration of transmission constraints. The results for this case are shown in Fig. 2(a) (a) and (b) and detailed generation dispatch (purple numbers) and demand values (green numbers) are shown in Fig. 2(b). Fig. 2(a) (a) shows results for Case I, with the black line showing the generation cost curve (with generators ordered from the cheapest to the most expensive) and the purple line showing the total carbon emission curve for Case I. The total carbon emissions $E_{tot}$ is calculated by multiplying the amount of generation with the corresponding carbon intensities, i.e., $E_{tot}=\sum_{g\in\mathcal{G}}e_{G,g}\cdot P_{G,g}$ based on the generation dispatch order. The red and purple points on the black generation cost curve represent the dispatch results for two cases, respectively. The purple shadowed areas highlight the effects of our model compared with the case without considering carbon emissions. Similarly, Fig. 2(a) (b) shows results for Case II, with the black line showing the generation cost curve (which is the same as for Case I) and the green line showing the total emission curve for Case II.
@@
Our first observation from Fig. 2(a), is that the only impact of considering the carbon-dependent utility is that the consumers are less willing to pay for electricity with a non-zero average carbon footprint and hence reduce their electricity demand. The generation merit order (represented by the generator cost curve) creates an ordering of the generators. The demand reduction sequentially impacts the generators, starting with the marginal one, which reduces production, and if needed, continues in descending order. If the affected generators have non-zero emissions, reductions in demand have the intended effect of decreasing emissions, as witnessed in both Case I and Case II. However, the fact that carbon-aware loads can only sequentially impact the generators according to the aforementioned ordering, which is not related to the ranking of generators by their emissions, is a fundamental limit of carbon-aware load shifting in our equilibrium model.
@@
Our second observation from Fig. 2(a) is that the demand reduction and associated reduction in emissions is very different between Case I and II. In Case I, the generation dispatch includes low-cost, carbon-intensive generation which leads to a high average carbon value $\lambda=0.913$. Since $\lambda$ is high, the carbon-dependent consumer utility is much lower than the normal consumer utility, leading to a large reduction in demand from 48 to 32 MW, or 33%, when considering carbon. However, since the marginal generator has low emissions, this reduction in demand only decreases emissions from 37 to 29.2 tons, or 21%. This suggests that each unit of demand reduction achieves a lower-than-average impact on emissions in Case I. Conversely, in Case II, the generation dispatch includes low-cost, low-carbon generation which leads to a low average carbon value $\lambda=0.4$. With a lower $\lambda$, the carbon-dependent consumer utility is closer to the normal consumer utility, and the demand is only reduced to 46.7 MW, or less than 3%, when considering carbon. This reduction in demand reduces emissions from 20 to 18.7 tons, or 6.5%. Thus, in Case II, each unit of demand reduction achieves a higher-than-average impact on emissions, but the overall demand and carbon reductions are small.
@@
(c) Results comparison for transmission-constrained cases.
2.2.2. Transmission-constrained Case
@@
To explain the impacts of transmission line constraints in our equilibrium model, we reduce the transmission capacity on the transmission line from Bus $2$ to Bus $3$ so it becomes congested. This congestion increases power generation at Bus 2 and decreases power generation at Bus 1. Fig. 2(c) shows the dispatch for the base case ($c_{D}=0$) and Case I and Case II ($c_{D}=20$), all with transmission line limits. For Case I, the shift of generation from generator 1 to 2 increases the average carbon emissions from $\lambda=0.608$ (for the electricity pool case) to $\lambda=0.767$ in the base transmission-constrained case (Case I in Fig. 3a). Introducing carbon-sensitive loads reduces the load from 48 to 32 MW, as in the case without transmission constraints. However, due to congestion, both generators 1 and 2 are now marginal and are impacted by the demand reduction, leading to a slightly higher reduction in emissions of 22.5%. Nevertheless, the average emissions increase to $\lambda=0.89$ after the load reduction. For Case II, shifting generation from generator 1 to 2 reduces the average carbon emissions to $\lambda=0.433$. This $\lambda$ is low enough to allow consumers to continue to consume their maximum load, and as a result, the generation dispatch does not change after we introduce carbon-sensitive loads. This is despite the fact that the two marginal generators in this case have higher emissions than in Case I.
@@
This example highlights that transmission constraints impact the marginal generators serving carbon-sensitive loads and thus impact the effectiveness of demand reductions from carbon-sensitive loads.
3. A Multi-Period Case with Temporal Load Shifting
Section 2 introduces an equilibrium model for a single time slot, in which consumers contribute to carbon reductions by curtailing their electricity demand from the maximum down to the minimum level when the carbon-dependent term $r_{D,d}-\lambda\cdot c_{D,d}$ is sufficiently small (specifically, when it falls below $p_{i:d\in\mathcal{D}_{i}}$). However, a more impactful and practical strategy for carbon reductions involves temporal load shifting, where consumers reallocate their fixed total electricity consumption across periods. To realistically model this strategy, it is crucial to also account for the time-varying nature of renewable generation. Thus, in this section, we formulate the equilibrium model in a multi-time-scale setting that considers both temporal load shifting and the dynamics of maximum renewable generation limits.
3.1. Model Formulation
Different from the single-period model in Section 2, each participant in the multi-period model needs to optimize their behaviors across considered periods under time-varying parameters, and detailed optimization problems are introduced below.
Generators: Given the electricity price at each time $p_{i,t}$ as input, the generator solves the following optimization problem to maximize their profit across $T$ time periods:
(6a) $\max_{P_{G,g,t}}\$$\sum_{t\in\mathcal{T}}(p_{i:g\in\mathcal{G}_{i},t}-c_{G,g,t})P_{G,g,t}$
(6b) $s.t.\$$P_{G,g}^{\min}\leq P_{G,g,t}\leq P_{G,g,t}^{\max},\ \forall t,$
where $\mathcal{T}=[1,2,\cdots,T]$. We assume that the maximum power $P_{G,g,t}^{\max}$ of all renewable generators are time-varying values, whereas the non-renewable values are fixed across time.
Consumers: Given time-varying electricity prices $p_{i,t}$ and average carbon signals $\lambda_{t}$ calculated by ISO in the later part, each consumer maximizes their utilities by solving the following problem:
(7a) $\max_{P_{D,d,t}}\$$\sum_{t\in\mathcal{T}}(r_{D,d,t}-p_{i:d\in\mathcal{D}_{i},t}-\lambda_{t}\cdot c_{D,d,t})\cdot P_{D,d,t}$
(7b) $s.t.\$$P_{D,d}^{\min}\leq P_{D,d,t}\leq P_{D,d}^{\max},\ \forall t,$
(7c) $\sum_{t\in\mathcal{T}}P_{D,d,t}=\mathcal{P}_{D,d}.$
where $\mathcal{P}_{D,d}$ is the fixed total electricity consumption across $T$ time slots for the consumer $d$. Note that in the multi-period case, the load schedule is dictated by the priority order of $r_{D,d,t}-p_{i:d\in\mathcal{D}_{i},t}-\lambda_{t}\cdot c_{D,d,t}$ across different time periods $t$. Specifically, consumption is prioritized for time slots with the larger values of $r_{D,d,t}-p_{i:d\in\mathcal{D}_{i},t}-\lambda_{t}\cdot c_{D,d,t}$.
Transmission Owners: The transmission owner aims to maximize their profits across $T$ time periods:
(8a) $\max_{\theta_{i,t}}\$$\sum_{t\in\mathcal{T}}\sum_{i\in\mathcal{N}}p_{i,t}\cdot\left(\sum_{j:(i,j)\in\mathcal{L}}\beta_{ij}(\theta_{j,t}-\theta_{i,t})\right)$
(8b) $s.t.\$$-F_{ij}^{\rm{lim}}\leq\beta_{ij}(\theta_{i,t}-\theta_{j,t})\leq F_{ij}^{\rm{lim}},\quad\forall(i,j)\in\mathcal{L},\ \forall t,$
(8c) $\theta_{ref,t}=0,\ \forall t.$
ISO: The ISO clears the market by determining electricity prices and provides average carbon signals. Specifically, they set electricity prices $p_{i,t}$ by ensuring the supply-demand balance constraint
(9) $\sum_{d\in\mathcal{D}_{i}}P_{D,d,t}+\sum_{j:(i,j)\in\mathcal{L}}\beta_{ij}(\theta_{i,t}-\theta_{j,t})=\sum_{g\in\mathcal{G}_{i}}P_{G,g,t},\ \forall i\in\mathcal{N},\ \forall t.$
Average carbon emissions can be calculated by:
(10) $\lambda_{t}\sum_{d\in\mathcal{D}}P_{D,d,t}=\sum_{g\in\mathcal{G}}e_{G,g,t}\cdot P_{G,g,t},\ \forall t.$
Both $p_{i,t}$ and $\lambda_{t}$ are parameters in the optimization problems (6)-(8).
Similar to the single-period model, the overall Nash equilibrium problem for the multi-period case consists of the collection of parametric optimization problems (6), (7), (8) and the parametric complementarity conditions (9) and (10).
3.2. Three-bus Example Illustration
To understand the impacts of average carbon signals on our multi-period equilibrium model with temporal load shifting, we consider the transmission-constrained three-bus example in Section 2. Specifically, we consider Case I in Table 1, under which the generator $2$ owns the highest generation cost but the lowest carbon factor of generating unit power, of which we assume the maximum power limit is time-varying. To make it simple, we only consider $T=3$ (longer time periods and more complex cases will be discussed for the IEEE RTS-GMLC system in numerical studies), $r_{D,d,t}$, $e_{G,g,t}$, and $c_{G,g,t}$ are fixed for each consumer across time, and the maximum generation values for generators 1 and 3 are fixed for all time slots at 2020MW and 1515MW, respectively, whereas the maximum power limit for generator 2 varies according to $P_{G,2,t}^{\max}=[5,15,30]$MW. The transmission capacity on the transmission line from Bus $1$ to Bus $2$ is set at $F_{12}^{lim}=8$MW (with the remaining two transmission lines unlimited) and the total consumption for each consumer across all periods is $\mathcal{P}_{D,d}=3*90\%*P_{D,d}^{\max}$.
Table 2. Impact of Temporal Load Shifting on Power Dispatch, Load Dispatch, Carbon Emissions, and Generation Costs.
Case definition Power dispatch Load dispatch Ave. Carbon Tot. Carbon Gen Cost Tot. Carbon Ave. Carbon
$g_{1,1}$ $g_{1,2}$ $g_{1,3}$ $g_{2,1}$ $g_{2,2}$ $g_{2,3}$ $g_{3,1}$ $g_{3,2}$ $g_{3,3}$ $d_{1,1}$ $d_{1,2}$ $d_{1,3}$ $d_{2,1}$ $d_{2,2}$ $d_{2,3}$ $d_{3,1}$ $d_{3,2}$ $d_{3,3}$ $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $t_{1}$ $t_{2}$ $t_{3}$
$c_{D,d}=0$ 19.5 19.9 18.7 3.9 11.5 11.1 15 15 15 5.9 5.4 4.9 17.1 23.7 24 15.4 17.3 15.9 0.72 0.63 0.64 27.6 29.2 28.7 999.6 85.2 0.66
$c_{D,d}=30$ 16.6 20 20 2 13 13 15 15 15 4.2 6 6 16.8 24 24 12.6 18 18 0.75 0.62 0.62 25.2 29.8 29.8 1002.9 84.6 0.65
Table 3.2 presents a comparative analysis of temporal load shifting between the base case $c_{D}=0$ and the case with $c_{D}=30$. Focusing on the consumer’s problem in (7), we have that given the same $r_{D,d,t}$ across all time, if the condition
$p_{i:d\in\mathcal{D}_{i},(1)}$ $+\lambda_{(1)}\cdot c_{D,d}\geq p_{i:d\in\mathcal{D}_{i},(2)}+\lambda_{(2)}\cdot c_{D,d}\geq$
(11) $\cdots\geq p_{i:d\in\mathcal{D}_{i},(T)}+\lambda_{(T)}\cdot c_{D,d}$
holds, then the corresponding power consumption follows
(12) $P_{D,d,(1)}\leq P_{D,d,(2)}\leq\cdots\leq P_{D,d,(T)}.$
This can be validated from the results in the case of $c_{D}=30$. The prices $p_{i,t}$ are all 10\$/MWh except for $p_{1,1}=8$\$/MWh and $p_{3,1}=9.1$\$/MWh under this case. The values of $p_{i:d\in\mathcal{D}_{i},t}+\lambda_{t}\cdot c_{D,d}$ at the latter two time slots are smaller than that at $t=1$. Hence, more electricity is consumed at the latter two time slots. Note that this load shifting is driven by lower carbon emissions at $t=2$ and $t=3$ (because of $p_{i,1}<p_{i,2}=p_{i,3}$ for consumers 1 and 3 and equal for consumer 2). This temporal load shifting aligns well with the trend of renewable power availability ($P_{G,2,t}^{\max}=[5,15,30]$).
A comparison between $c_{D}=0$ and $c_{D}=30$ reveals that the influence of average carbon signals only marginally increases renewable generation (from Generator 2) from 26.5MWh to 28MWh. Consequently, total carbon emissions experience a slight decline from 85.2tCO2 to 84.6tCO2, accompanied by a cost increase from \$999.6 to \$1002.9. Another interesting observation is that although average carbon emissions decrease at $t=2$ and 3, total emissions at these times increase due to higher consumption levels happening at the same time for $c_{D}=30$ (compared to $c_{D}=0$). Furthermore, in both cases, the generation dispatch results obtained by solving Problem (6) follow the generation cost merit order, with lower-cost generators being dispatched first. For instance, Generator 3, having the lowest cost, operates at its maximum output of 15 MWh across all time periods, while the total power output of Generator 1 consistently exceeds that of Generator 2, which has the highest cost.
This example highlights again why the average carbon signal may not be the best choice, as it promotes peak consumption occurring simultaneously, i.e., incentivizing increased consumption during periods of lower average carbon emissions. While this may reduce the average carbon intensity at certain times, it can lead to an overall increase in total carbon emissions during those periods (as shown in the above example). Furthermore, as noted in (Ruiz and Rudkevich, 2010), depending on the specific power system conditions, certain consumers may need to take actions contrary to others in order to effectively reduce system-wide total carbon emissions. It is evident that average carbon signals lack the granularity needed to facilitate such coordinated and differentiated responses.
4. Numerical Studies
@@
This section conducts extensive numerical studies to demonstrate how the equilibrium method could impact electricity market clearing. All the tests are performed on a laptop computer with an Apple M3 Pro CPU and 36GB RAM. Our equilibrium problem is reformulated as a system of constraints formed by the KKT conditions of Problems  (1) – (3), together with the complementarity constraints  (4) and  (5). This system is then implemented in GAMS (or GAMSpy) (GAMs, 2024) using the Extended Mathematical Programming (EMP) framework (Kim and Ferris, 2019) and solved numerically using the PATH solver (Dirkse and Ferris, 1995). We adopt the IEEE RTS-GMLC system (Barrows et al., 2019) for numerical analysis. This system has $73$ buses, $158$ generators, and $120$ lines. We assign emission factors for each generator based on their assigned fuel type and data from the US Department of Energy (Office of Energy Policy and Systems Analysis, 2016). Specifically, we assign $e_{G}=\{0.6042,0.7434,0.9606\}$ for natural gas, oil, and coal generators, respectively, and assume that solar, wind, and hydro have $e_{G}=0$. For simplicity, we consider each non-zero load in the system as a consumer, leading to a total of $51$ consumers located at different buses. We draw consumer utility values $u_{D}$ from a uniform distribution in the range of $[20,80]$ \$/MWh. According to (World Bank Group, 2023), carbon prices in the worldwide emission trading markets range from $10$ to 100100\$/tons. Based on these statistics, we assume that consumers will have carbon prices $c_{D}$ in the range between $[0,100]$ in the following experiments. For consumers’ demand flexibility, we use the values in the IEEE RTS-GMLC system as the maximum values and 80% of these values as the minimum values.
4.1. Evaluation Metrics and Benchmark Formulations
@@
In our case study, we consider the following quantities when evaluating market-clearing solutions:
@@
We consider five benchmark models for our analysis:
@@
  • Market Clearing with Demand Flexibility (MCDF): In this case, consumers have flexibility in their electricity consumption but do not consider carbon costs. Note that this model is equivalent to a version of our equilibrium model where all consumers set carbon costs $c_{D}=0$.
  • Market Clearing with Minimizing Carbon Emissions (MCE): For comparison, we also involve a case to minimize carbon emissions of the whole system, where constraints are the same as MCDF, but the objective function is to minimize the total carbon emission as follows:
    (13) $\min_{P_{G},P_{D},\theta}\left(\sum_{g\in\mathcal{G}}e_{G,g}P_{G,g}\right).$
  • Carbon Tax Method (CTM): One common effort to reduce carbon emissions is adding carbon taxes on electric generators (Baranzini et al., 2000; Fischer and Newell, 2008). Mathematically, the problem can be detailed as follows:
    $\max_{P_{G},P_{D},\theta}$ $\left(\sum_{d\in\mathcal{D}}r_{D,d}P_{D,d}-\sum_{g\in\mathcal{G}}\left(c_{G,g}+C_{co_{2}}e_{G,g}\right)P_{G,g}\right)$
[…]
    where $C_{co_{2}}\in\mathbb{R}$ is the carbon tax, which is usually determined by energy-related governments. From a game-theoretic perspective, this carbon tax model is also equivalent to an equilibrium model, where each generator maximizes their profits considering both price and carbon emissions in their decisions. We defer the detailed equilibrium formulation for the carbon tax method to Appendix B.
  • Sequential Method (SM): We consider the sequential method first developed to represent geographical load shifting in (Lindberg et al., 2021b) and extended to consider multiple time steps in (Gorka et al., 2025). This method is representative of current practice and consists of three steps:
@@
    Step $3$: Finally, we solve another market clearing similar to the MCFD, representing the intra-day electricity market, with updated load profiles. We denote the average carbon emission based on these market clearing results as $\lambda_{a}$.
    One critical issue with this sequential method is that there is a time lag in the signal $\lambda_{e}$ received by the consumer. Specifically, consumers only receive the carbon emission signal a posteriori, after the market is cleared, and do not know the impacts of their behaviors on the next market clearing results. This is in contrast to our equilibrium formulation, which allows consumers to possess “perfect” knowledge of carbon emission intensity signals.
4.2. Analysis for Different Carbon Costs
We first analyze the impact of different carbon costs on the equilibrium results, assuming that all consumers are willing to submit non-zero carbon costs to help reduce carbon emissions. We compare cases where the carbon costs of consumers are randomly drawn from a uniform distribution with ranges $[10,40]$, $[30,60]$, $[50,80]$, and $[80,100]$, respectively. Table 3 shows the results. The top rows represent the two carbon-agnostic formulations (MCFD and MCDF), and the average carbon cost increases as we go further down in the rows.
Table 3. The impact of different carbon costs on generation dispatch, system emissions, and other components in equilibrium models.